OpenStudy (anonymous):

evaluate:∫2dy/(3y-4).. need the solution and answer plz help me to solve this problem,,,tnx,,

OpenStudy (anonymous):

why?

OpenStudy (zarkon):

\[u=3y-4\]

OpenStudy (anonymous):

∫2dy/(3y-4) = 2∫dy/(3y-4) t = 3y-4 => dt = 3dy => dy = dt/3 2∫dy/(3y-4) = 2∫dt/3t = (2/3)*ln t + C ∫2dy/(3y-4) =[ln (3y-4)^2]/3 + C

OpenStudy (amistre64):

it has to do with reading the future; we need to recognize where 1/(3y-4) comes from derivatively. ln(3y-4) derives down to 3/(3y-4) times some constant which was pulled out to begin with

OpenStudy (amistre64):

\[\frac{2}{3y-4}=\frac{3}{3y-4}*N\] \[\frac{2}{3y-4}*\frac{3y-4}{3}=N\] \[\frac{2}{3}=N\] \[\frac{2}{3}ln(3y-4)\text{ is at least a reasonable result}\]

OpenStudy (amistre64):

the problems are designed to get you to think "outside" the linear box and get you used to thinking ahead of the problem ...

OpenStudy (zarkon):

the absolute value should really be included in the answer unless you write the final solution the way giorgia wrote it.